OpenStudy (j2lie):

How do you solve mixture problems without using a table? @agent0smith

OpenStudy (johnweldon1993):

What's an example?

OpenStudy (j2lie):

?

OpenStudy (j2lie):

How many liters of a 70% alcohol solution must be added to 50 liters of a 40% alcohol solution to produce a 50% alcohol solution? This is a example of a mixture problem.

OpenStudy (agent0smith):

write an equation, x is the #liters of 70%, y is the #liters of 40% 0.7x + 0.4(50) = 0.5(50+x)

OpenStudy (j2lie):

Thank you.

OpenStudy (j2lie):

how do you solve that?

OpenStudy (agent0smith):

Try it first.

OpenStudy (j2lie):

I did it but I got a negative.

OpenStudy (agent0smith):

Show your work.

OpenStudy (j2lie):

0.7x+0.4(50)=o.5(50+x) 7x+4(500)=5(500+x) 7x+2000=2500+5x -5x-2500 2x+-500 x=-250%

OpenStudy (agent0smith):

0.7x+0.4(50)=o.5(50+x) 7x+4(500)=5(500+x) explain what happened here. It seems like you wanted to multiply by 10, but you only multiplied SOME terms by 10. You don't need to do that anyway.

OpenStudy (j2lie):

I multiplied all by 10.

OpenStudy (j2lie):

Oh wait I forgot to multiply x by 10.

OpenStudy (agent0smith):

You don't need to multiply by 10.

OpenStudy (j2lie):

how do you do it without multiplying by 10? And stop being grumpy please.

OpenStudy (agent0smith):

This is not being grumpy lol, you just do it without multiplying by 10. 0.7x + 0.4(50) = 0.5(50+x) distribute... 0.7x + 20 = 25 + 0.5x try from now

OpenStudy (j2lie):

If you're not smiling then you're grumpy.

OpenStudy (agent0smith):

lol okay. If you can deal with simple decimals like those there's no need to multiply by 10.

OpenStudy (j2lie):

If I can't deal with a simple decimal, how am I suppose to deal with two decimals.

OpenStudy (agent0smith):

Why can't you deal with simple decimals like 0.7? You seem fine with algebra

OpenStudy (j2lie):

I'm not a smart girl.

OpenStudy (agent0smith):

0.7x + 20 = 25 + 0.5x get all the x's on one side by subtracting 0.5x 0.7x - 0.5x + 20 = 25 subtract 20 on both sides 0.7x - 0.5x = 5 0.7x - 0.5x is 0.2x, so 0.2x = 5

OpenStudy (j2lie):

x=2.5

OpenStudy (agent0smith):

0.2x = 5 divide both sides by 0.2 x = 105

OpenStudy (j2lie):

how come it is 105?